40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5235     Entry time: Mon Aug 15 10:02:27 2011
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: BS OSEM adjustment done 

[Suresh / Kiwamu]

 Adjustment of the OSEMs on BS has been done.

All the bad suspensions (#5176) has been adjusted. They are waiting for the matrix inversion test.

ELOG V3.1.3-